# Amortized Sampling with Transferable Normalizing Flows

### v3

@@ Abstract
Efficient equilibrium sampling of molecular conformations remains a core challenge in computational chemistry and statistical inference. Classical approaches such as molecular dynamics or Markov chain Monte Carlo inherently lack amortization; the computational cost of sampling must be paid in full for each system of interest. The widespread success of generative models has inspired interest towards overcoming this limitation through learning sampling algorithms. Despite performing competitively with conventional methods when trained on a single system, learned samplers have so far demonstrated limited ability to transfer across systems. We demonstrate that deep learning enables the design of scalable and transferable samplers by introducing Prose, a 285 million parameter all-atom transferable normalizing flow trained on a corpus of peptide molecular dynamics trajectories up to 8 residues in length. Prose draws zero-shot uncorrelated proposal samples for arbitrary peptide systems, achieving the previously intractable transferability across sequence length, whilst retaining the efficient likelihood evaluation of normalizing flows. Through extensive empirical evaluation we demonstrate the efficacy of Prose as a proposal for a variety of sampling algorithms, finding a simple importance sampling-based finetuning procedure to achieve competitive performance to established methods such as sequential Monte Carlo. We open-source the Prose codebase, model weights, and training dataset, to further stimulate research into amortized sampling methods and finetuning objectives.
@@ 1 Introduction
  • We study the use of Prose as a proposal distribution for different Monte Carlo algorithms, finding the learned proposal to be sufficiently powerful for accurate sampling with standard SNIS, which does not require tuning of sampling parameters. Furthermore, resampled generations can be used for efficient finetuning of Prose on previously unseen systems.
@@ 3.1 Architecture of Prose
In the image setting, Zhai et al. (2024) employ only an identity and flip permutation to the sequence of image tokens. Similarly, when applying TarFlow to peptide systems Tan et al. (2025) employ only an identity and flip permutation on the ordering defined by the PDB format, in which the sequence is defined per-residue starting with backbone atoms followed by sidechain atoms. Whilst a simple identity and flip may be appropriate for the regular grid of image data, we argue this to be suboptimal for the diversity of geometric interactions present in molecular systems. This motivates our introduction of chemistry-aware sequence permutations, defined to promote effective peptide modeling. We define the backbone permutation, such that the backbone atoms $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=0}^{L-1}$ (with associated hydrogens) for all residues are located at the start of the sequence, and followed by the sidechains. By processing the coordinates of the backbone atoms at the start of the sequence, the model refines the global structure of the peptide as a contiguous sequence. Crucially, the sidechain positions are subsequently able to causally attend to the full backbone structure, hence enabling local updates to be influenced by global structure. We further employ a backbone-flip permutation to provide additional diversity to the autoregressive modeling.
@@ 4.3 Scale transferability of Prose
We present metrics for Prose and baseline methods in Table˜2. Prose achieves the strongest performance on all metrics aside from dipeptide $\mathcal{T}\text{‑}\mathcal{W}_{2}$, where it is marginally outperformed by TarFlow, confirming it to be a strong SNIS proposal for peptide systems of varying sequence length. ECNF++ performs very poorly on $E\text{‑}\mathcal{W}_{2}$ on both dipeptides and tetrapeptides, seemingly unable to learn an effective vector field when trained on tetrapeptides. TimeWarp is the strongest non-Boltzmann generator baseline, with both BioEmu and UniSim attaining high values of $E\text{‑}\mathcal{W}_{2}$. However, we note the training data for these pretrained models does not correspond exactly with our evaluation data and hence they are not directly comparable. Fig.˜1 further confirms the success of Prose with SNIS as an amortized sampler, surpassing the performance of a baseline MD trajectory on the critical $\mathcal{T}\text{‑}\mathcal{W}_{2}$ and $\mathrm{TICA}\text{‑}\mathcal{W}_{2}$ describing metastable state coverage w.r.t. both energy evaluations and GPU walltime. We additionally present qualitative results on the unseen octopeptide DGVAHALS in Fig.˜3, demonstrating the unprecedented scalability of Prose; further results are provided in Appendix˜F.
Figure 3: Prose accurately samples from the Boltzmann distributions of unseen octopepitde system. Empirical results for sampling from DGVAHALS peptide system, not present in training data. Energy histogram (left) for reference MD data, Prose proposal and Prose reweighted using SNIS, demonstrate fine-grained detail accuracy. TICA plots for MD (center) and SNIS-reweighted Prose (right) illustrate mode coverage.
@@ 4.5 Sampling algorithms
Having established the unmatched performance of Prose in the standard Boltzmann generator framework, we now consider alternative sampling algorithms made tractable by its efficient likelihood. We evaluate SNIS, SMC in continuous time, SMC in discrete time, and the simple instantiation of self-improvement defined in Section˜3.2. All methods are permitted a budget of $10^{6}$ energy evaluations, further details on method configurations are provided in Appendix˜E. Metric results are presented in Table˜4. These results reveal the surprising result that, given a suitably strong proposal distribution, SNIS is competitive with both SMC variants, despite requiring no tuning. While SMC discrete achieves the best value of $E\text{‑}\mathcal{W}_{2}$ on octopeptides, both SMC variants introduce a notable deterioration of macrostructure metrics at this scale when compared to SNIS. Furthermore, the performance of SNIS with self-improvement at improving the tetra- and octopeptide $E\text{‑}\mathcal{W}_{2}$ provides strong evidence in favor of proposal finetuning within sampling methods, as an alternative to resource allocation solely on annealing-based methods. We provide results for the unseen RLMM system in Fig.˜4, illustrating the superior mode coverage of Prose with SNIS over a MD baseline given an allocation of $10^{6}$ energy evaluations, further evidence of successful amortized sampling with Prose.
@@ 5 Related Work
Boltzmann generators are machine learning-based samplers that train likelihood-based models and employ inference-time SNIS to achieve consistent sampling of the target density (Noé et al., 2019). A major limitation of standard Boltzmann generators is the need to train the proposal model on a dataset of true density samples, motivating methods that transfer between molecular systems. Accordingly, Jing et al. (2022) propose a transferable Boltzmann generator operating on the torsion angles of small molecules, and Klein and Noe (2024) develop a Boltzmann generator operating on Cartesian coordinates that transfers between dipeptide systems. However, the scalability of such methods has remained limited due to the difficulty of designing expressive generative models that possess efficient and accurate likelihood evaluations. In particular the use of continuous normalizing flows implies a large cost to proposal likliehoods due to the need to integrate the vector field divergence (Grathwohl et al., 2019). Schopmans and Friederich (2025) replace the single-step SNIS of standard Boltzmann generators with a temperature-annealing sequence of normalizing flows, performing SNIS with each flow to sample from a given target distribution.
@@
The widespread empirical success of generative modeling has inspired many approaches to machine learning-based sampling. Boltzmann emulators, like Boltzmann generators, seek uncorrelated sampling of the target density, but forgo efficient likelihood evaluation in favor of scalable generative modeling on large pre-collected datasets (Abdin and Kim, 2024; Wayment-Steele et al., 2024; Lewis et al., 2025). In this case the lack of efficient likelihood evaluation precludes the use of Monte Carlo estimators such as SNIS. Diffusion samplers, propose novel objectives for training diffusion models in the absence of an empirical data distribution, both simulation-based (Berner et al., 2024; Vargas et al., 2023; Richter et al., 2024; Zhang and Chen, 2022; Vargas et al., 2024) and simulation-free (Akhound-Sadegh et al., 2024; Huang et al., 2021; De Bortoli et al., 2024). Notably, Havens et al. (2025) develop a diffusion sampler that is transferable across small molecules. Time coarseners are another family of model, in which ML is used to predlct large time transitions for simulation (Schreiner et al., 2023; Fu et al., 2023; Klein et al., 2023b; Daigavane et al., 2024; Yu et al., 2025), whereas methods like MDGen apply generative modeling to both the spatial and temporal dimensions of MD data (Jing et al., 2024). Lastly, several works integrate normalizing flows with classical Monte Carlo methods (Albergo et al., 2019; Arbel et al., 2021; Gabrié et al., 2021; Matthews et al., 2022; Midgley et al., 2023b; Hagemann et al., 2023).
@@ E.3 Sampling algorithm configurations
For both variants we generate a proposal set of $10^{4}$ samples. For the discrete variant, we perform 50 annealing steps, requiring two energy evaluations per step: one to update the samples and one for the Metropolis–Hastings correction. For the continuous variants, we perform 100 annealing steps. In both cases, resampling is performed at every step. For the continuous variant, Langevin dynamics is used with a step size $\sigma_{t}$ of $10^{-7}$ for dipeptides and tetrapeptides, and $\sigma_{t}$ of $10^{-8}$ for octopeptides. Details of the formulation are found in Section˜D.2. For the discrete variant, we apply Langevin dynamics with a step size of $10^{-5}$, followed by a Metropolis-Hastings step to accept or reject proposals. The step size is adaptively updated to maintain an acceptance rate of approximately 60%, under the assumption of sufficient smoothness in the intermediate densities. Further details of discrete-time AIS are found in Section˜D.3.
@@ F.3 Dataset scaling
Figure 9: Wasserstein-distance metrics for Prose trained on variants of ManyPeptidesMD. Upper row: ManyPeptidesMD contains $200\text{\,}\mathrm{ns}$ trajectories for all training sequences, we train on a variant with $50\text{\,}\mathrm{ns}$ and $100\text{\,}\mathrm{ns}$ respectively. Lower: ManyPeptidesMD contains 21,700 sequences, we train using 25% and 50% of the total. Evaluation metrics computed using SNIS on 30 octopeptide sequences.
@@ F.5 JSD evaluation
bsectionEffective sample per second
@@
F.6 Octopeptide Ramachandran plots
Figure 10: Ramachandran plots for DGVAHALS unseen octopeptide system. Reference molecular dynamics (left column), Prose proposal (center column), Prose SNIS with $10^{5}$ samples (right column).
F.7 Temperature Plops

### v4

@@ Abstract
Efficient equilibrium sampling of molecular conformations remains a core challenge in computational chemistry and statistical inference. Classical approaches such as molecular dynamics or Markov chain Monte Carlo inherently lack amortization; the computational cost of sampling must be paid in full for each system of interest. The widespread success of generative models has inspired interest towards overcoming this limitation through learning sampling algorithms. Despite performing competitively with conventional methods when trained on a single system, learned samplers have so far demonstrated limited ability to transfer across systems. We demonstrate that deep learning enables the design of scalable and transferable samplers by introducing Prose, a 285 million parameter all-atom transferable normalizing flow trained on a corpus of peptide molecular dynamics trajectories up to 8 residues in length. Prose draws zero-shot uncorrelated proposal samples for arbitrary peptide systems, achieving the previously intractable transferability across sequence length, whilst retaining the efficient likelihood evaluation of normalizing flows. Through extensive empirical evaluation we demonstrate the efficacy of Prose as a proposal for a variety of sampling algorithms, finding a simple importance sampling-based fine-tuning procedure to achieve competitive performance to established methods such as sequential Monte Carlo. We open-source the Prose codebase, model weights, and training dataset, to further stimulate research into amortized sampling methods and objectives.
@@ 1 Introduction
  • We study the use of Prose as a proposal distribution for different Monte Carlo algorithms, finding the learned proposal to be sufficiently powerful for accurate sampling with standard SNIS, which does not require tuning of sampling parameters. Furthermore, resampled generations can be used for efficient fine-tuning of Prose on previously unseen systems.
@@ 3.1 Architecture of Prose
In the image setting, Zhai et al. (2024) employ only an identity and flip permutation to the sequence of image tokens. Similarly, when applying TarFlow to peptide systems Tan et al. (2025) employ only an identity and flip permutation on the ordering defined per-residue starting with backbone atoms followed by sidechain atoms. Whilst a simple identity and flip may be appropriate for the regular grid of image data, we argue this to be suboptimal for the diversity of geometric interactions present in molecular systems. This motivates our introduction of chemistry-aware sequence permutations, defined to promote effective peptide modeling. We define the backbone permutation, such that the backbone atoms $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=0}^{L-1}$ (with associated hydrogens) for all residues are located at the start of the sequence, and followed by the sidechains. By processing the coordinates of the backbone atoms at the start of the sequence, the model refines the global structure of the peptide as a contiguous sequence. Crucially, the sidechain positions are subsequently able to causally attend to the full backbone structure, hence enabling local updates to be influenced by global structure. We further employ a backbone-flip permutation to provide additional diversity to the autoregressive modeling.
@@ 4.3 Scale transferability of Prose
We present metrics for Prose and baseline methods in Table˜2. Prose achieves the strongest performance on all metrics aside from dipeptide $\mathcal{T}\text{‑}\mathcal{W}_{2}$, where it is marginally outperformed by TarFlow, confirming it to be a strong SNIS proposal for peptide systems of varying sequence length. ECNF++ performs very poorly on $E\text{‑}\mathcal{W}_{2}$ on both dipeptides and tetrapeptides, seemingly unable to learn an effective vector field when trained on tetrapeptides. TimeWarp is the strongest non-Boltzmann generator baseline, with both BioEmu and UniSim attaining high values of $E\text{‑}\mathcal{W}_{2}$. However, we note the training data for these pretrained models does not correspond exactly with our evaluation data and hence they are not directly comparable. Fig.˜1 further confirms the success of Prose with SNIS as an amortized sampler, surpassing the performance of a baseline MD trajectory on the critical $\mathcal{T}\text{‑}\mathcal{W}_{2}$ and $\mathrm{TICA}\text{‑}\mathcal{W}_{2}$ describing metastable state coverage w.r.t. both energy evaluations and GPU walltime. We additionally present qualitative results on the unseen octapeptide DGVAHALS in Fig.˜3, demonstrating the unprecedented scalability of Prose; further results are provided in Appendix˜F.
Figure 3: Prose accurately samples from the Boltzmann distributions of unseen octapepitde system. Empirical results for sampling from DGVAHALS peptide system, not present in training data. Energy histogram (left) for reference MD data, Prose proposal and Prose reweighted using SNIS, demonstrate fine-grained detail accuracy. TICA plots for MD (center) and SNIS-reweighted Prose (right) illustrate mode coverage.
@@ 4.5 Sampling algorithms
Having established the unmatched performance of Prose in the standard Boltzmann generator framework, we now consider alternative sampling algorithms made tractable by its efficient likelihood. We evaluate SNIS, SMC in continuous time, SMC in discrete time, and the simple instantiation of self-improvement defined in Section˜3.2. All methods are permitted a budget of $10^{6}$ energy evaluations, further details on method configurations are provided in Appendix˜E. Metric results are presented in Table˜4. These results reveal the surprising result that, given a suitably strong proposal distribution, SNIS is competitive with both SMC variants, despite requiring no tuning. While SMC discrete achieves the best value of $E\text{‑}\mathcal{W}_{2}$ on octapeptides, both SMC variants introduce a notable deterioration of macrostructure metrics at this scale when compared to SNIS. Furthermore, the performance of SNIS with self-improvement at improving the tetra- and octapeptide $E\text{‑}\mathcal{W}_{2}$ provides strong evidence in favor of proposal fine-tuning within sampling methods, as an alternative to resource allocation solely on annealing-based methods. We provide results for the unseen RLMM system in Fig.˜4, illustrating the superior mode coverage of Prose with SNIS over a MD baseline given an allocation of $10^{6}$ energy evaluations, further evidence of successful amortized sampling with Prose.
@@ 5 Related Work
Boltzmann generators are machine learning-based samplers that train likelihood-based models and employ inference-time SNIS to achieve consistent sampling of the target density (Noé et al., 2019). A major limitation of standard Boltzmann generators is the need to train the proposal model on a dataset of true density samples, motivating methods that transfer between molecular systems. Accordingly, Jing et al. (2022) propose a transferable Boltzmann generator operating on the torsion angles of small molecules, and Klein and Noe (2024) develop a Boltzmann generator operating on Cartesian coordinates that transfers between dipeptide systems. However, the scalability of such methods has remained limited due to the difficulty of designing expressive generative models that possess efficient and accurate likelihood evaluations. In particular the use of continuous normalizing flows implies a large cost to proposal likelihoods due to the need to integrate the vector field divergence (Grathwohl et al., 2019). Schopmans and Friederich (2025) replace the single-step SNIS of standard Boltzmann generators with a temperature-annealing sequence of normalizing flows, performing SNIS with each flow to sample from a given target distribution.
@@
The widespread empirical success of generative modeling has inspired many approaches to machine learning-based sampling. Boltzmann emulators, like Boltzmann generators, seek uncorrelated sampling of the target density, but forgo efficient likelihood evaluation in favor of scalable generative modeling on large pre-collected datasets (Abdin and Kim, 2024; Wayment-Steele et al., 2024; Lewis et al., 2025). In this case the lack of efficient likelihood evaluation precludes the use of Monte Carlo estimators such as SNIS. Diffusion samplers, propose novel objectives for training diffusion models in the absence of an empirical data distribution, both simulation-based (Berner et al., 2024; Vargas et al., 2023; Richter et al., 2024; Zhang and Chen, 2022; Vargas et al., 2024) and simulation-free (Akhound-Sadegh et al., 2024; Huang et al., 2021; De Bortoli et al., 2024). Notably, Havens et al. (2025) develop a diffusion sampler that is transferable across small molecules. Time coarseners are another family of model, in which ML is used to predict large time transitions for simulation (Schreiner et al., 2023; Fu et al., 2023; Klein et al., 2023b; Daigavane et al., 2024; Yu et al., 2025), whereas methods like MDGen apply generative modeling to both the spatial and temporal dimensions of MD data (Jing et al., 2024). Lastly, several works integrate normalizing flows with classical Monte Carlo methods (Albergo et al., 2019; Arbel et al., 2021; Gabrié et al., 2021; Matthews et al., 2022; Midgley et al., 2023b; Hagemann et al., 2023).
@@ E.3 Sampling algorithm configurations
For both variants we generate a proposal set of $10^{4}$ samples. For the discrete variant, we perform 50 annealing steps, requiring two energy evaluations per step: one to update the samples and one for the Metropolis–Hastings correction. For the continuous variants, we perform 100 annealing steps. In both cases, resampling is performed at every step. For the continuous variant, Langevin dynamics is used with a step size $\sigma_{t}$ of $10^{-7}$ for dipeptides and tetrapeptides, and $\sigma_{t}$ of $10^{-8}$ for octapeptides. Details of the formulation are found in Section˜D.2. For the discrete variant, we apply Langevin dynamics with a step size of $10^{-5}$, followed by a Metropolis-Hastings step to accept or reject proposals. The step size is adaptively updated to maintain an acceptance rate of approximately 60%, under the assumption of sufficient smoothness in the intermediate densities. Further details of discrete-time AIS are found in Section˜D.3.
@@ F.3 Dataset scaling
Figure 9: Wasserstein-distance metrics for Prose trained on variants of ManyPeptidesMD. Upper row: ManyPeptidesMD contains $200\text{\,}\mathrm{ns}$ trajectories for all training sequences, we train on a variant with $50\text{\,}\mathrm{ns}$ and $100\text{\,}\mathrm{ns}$ respectively. Lower: ManyPeptidesMD contains 21,700 sequences, we train using 25% and 50% of the total. Evaluation metrics computed using SNIS on 30 octapeptide sequences.
@@ F.5 JSD evaluation
F.6 Effective sample per second
@@
F.7 Octapeptide Ramachandran plots
Figure 10: Ramachandran plots for DGVAHALS unseen octapeptide system. Reference molecular dynamics (left column), Prose proposal (center column), Prose SNIS with $10^{5}$ samples (right column).
F.8 Temperature plots